# PEVLM: Parallel Encoding for Vision-Language Models

### v2

@@ Results
The experiments show that PEVLM achieves an accuracy comparable to, and in some cases even surpassing, that of Full-Attention across all three evaluated models. Both APE and Star-Attention also approach or exceed the accuracy of Full-Attention when Full-Attention’s position encoding is retained. However, Block-Attention exhibits a notable accuracy drop compared to Full-Attention in most settings, primarily due to the lack of alignment-specific fine-tuning.
Although parallel encoding is generally expected to degrade performance, it can instead improve accuracy in long-context scenarios. This is because parallel encoding effectively shortens the total token length involved in each softmax computation, thereby mitigating the softmax feature degradation that occurs with long contexts (Veličković et al. 2025). As a result, the softmax output more robustly approximates an ideal sharp function, which ultimately enhances the performance of some models under long-context conditions.
@@
In this section, we evaluate the performance of PEVLM. To facilitate experimentation, we built PEVLM basied on the SGLang, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment.
Figure 4: We evaluate PEVLM under different block_size settings. The sink_size is set to (system prompt+one frame). "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs. "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
@@
Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights. Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model. The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM. All experiments were carried out on an NVIDIA H20-96G GPGPU. Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
Figure 5: We evaluate PEVLM under different sink_size settings. The block_size is set to (16 frames). "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs. "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
@@
As illustrated in Figure 4 and Figure 5, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×. Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected. As the block size or sink size increases, the acceleration effect of the attention module decreases significantly. This observation is consistent with the Equation 11. Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
@@
One key challenge in deploying VLMs on edge devices is the limited computational resources and strict latency requirements. To evaluate the acceleration performance of PEVLM under this scenario, we designed a simulation accordingly. A simplified experimental setup was implemented to assess the efficiency of PEVLM in edge deployment settings.
@@
Setup
We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure. The experiments are conducted on the H20-96B platform. We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare Full-Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct).
@@
As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s. PEVLM maintain relatively high accuracy even at 20s. Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
@@
Position Encoding
We further analyze the component with the greatest impact on accuracy: the use of sequential position embeddings. As discussed before, adopting sequential position embeddings instead of reusing position embeddings across blocks produces attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s expected behavior. As shown in Table 4, the use of sequential position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA. In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution (as shown in Figure 9 in the Appendix) remains relatively stable even when using parallel encoding.
@@
(128) PEVLM 58.34% 59.01%
Sink & Block Size
As analyzed in the third section, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 5.
Furthermore, we investigate the impact of sink size on model accuracy. As observed in the second section, it is essential to involve the initial frames of the video in the sink. This observation is validated in Figure 6, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
@@
Moreover, as shown in Figure 6, the accuracy of all models shows an upward trend as the sink size increases. However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance. Therefore, it may be necessary to conduct a simple test prior to deployment to determine an appropriate sink size. For the three models shown in Figure 6, setting the sink size between 14 and 18 already yields strong accuracy. If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy. For example, as demonstrated in the results of Table 2, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
@@
As shown in Figure 4, the inference latency of PEVLM also increases quadratically with block size, which is consistent with the experimental results shown in Figure 4. However, the block size has a more significant impact on performance because the coefficient of the block size’s quadratic term in Equation 11 is much larger than that of the sink size.
We investigate the impact of block size on model accuracy here. As shown in Figure 7, the model accuracy exhibits an upward trend as the block size increases, similar to sink size. However, due to the quadratic growth in latency, a trade-off arises between computational efficiency and accuracy.
@@ Conclusion
This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation. To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing. PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing sequential position embeddings. Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
@@
Appendix A Benchmark Details
@@
Appendix B Attention Weights Distribution of LLaVA-Video and LongVILA

### v3

@@ Results
As shown in Table 1, for shorter video scenarios, all methods achieve accuracy comparable to or even better than Full-Attention. Although parallel encoding is generally expected to degrade performance, it can actually improve accuracy in long-context scenarios. This is because the softmax function in attention mechanisms struggles to produce reliable outputs when the input length exceeds the training horizon (Veličković et al. 2025). However, parallel encoding effectively reduces the number of tokens involved in each softmax computation, thus alleviating the degradation of softmax behavior in long-context settings. As a result, the softmax output more robustly approximates an ideal sharp distribution, ultimately enhancing the performance of certain models under extended context lengths.
While all parallel encoding methods perform well on short videos, their effectiveness declines as the temporal length of the video increases, especially when video durations extend to several or even tens of minutes. In such scenarios, PEVLM consistently outperforms other parallel encoding approaches, particularly on the Qwen2.5-VL model. Qwen2.5-VL uses 3D-MROPE and is thus more sensitive to temporal order in videos. As APE and Star-Attention reuse position IDs across different context blocks, which severely distort temporal features, resulting in meaningless or garbled output. In contrast, Block-Attention reencodes positional information, which improves its accuracy on Qwen2.5-VL. PEVLM further improves accuracy by more rigorously preserving sequential positional encodings and incorporating a shared sink block, leading to better results. For LongVILA and LLaVA-Video, which use standard RoPE, the aforementioned issues are less pronounced. Nonetheless, PEVLM still outperforms other parallel encoding methods in long-video scenarios.
@@
In this section, we evaluate the performance of PEVLM.
Figure 4: PEVLM under different context_block_size settings. "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs. "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance. The sink_size is set to (system prompt+one frame).
Figure 5: PEVLM under different sink_size settings. "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs. "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance. The context_block_size is set to (16 frames).
@@
To facilitate experimentation, we built PEVLM based on the SGLang, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment. We conducted computational efficiency evaluations on the Qwen2.5-VL model. The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM. All experiments were carried out on an NVIDIA H20-96G GPGPU. Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
@@
As shown in Figure 4 and Figure 5, the end-to-end inference speed of the LLM improves by 2.58× to 3.14× as the block size and sink size vary. The attention module, in particular, achieves substantial acceleration, with speedups ranging from 7.79× to 33.81×, while the runtime of non-attention components remains largely stable. As the size of the sink block or context blocks increases, the speedup of the attention module correspondingly decreases. This trend is consistent with Equation 10, which indicates that the computational cost of attention grows quadratically with the size of the sink and context blocks. However, since non-attention components account for a considerable portion of total runtime, the overall end-to-end speedup degrades only moderately. However, as the context blocks size and sink block size increase to 16 frames, the proportion of attention in the total runtime increases from 6.4% to 22.9%, indicating that further increases in block size will lead to a more pronounced drop in overall performance. We will analyze the trade-off between performance degradation and potential accuracy gains from larger sink/context blocks size in the following sections.
@@
Deploying VLMs on edge devices presents significant challenges due to constrained computational resources and stringent latency requirements. To evaluate the practical benefits of PEVLM in such settings, we designed a latency-aware simulation that mimics real-world deployment conditions on resource-limited devices.
Setup
We adopt LongVideoBench as the evaluation benchmark and introduce a latency threshold during inference: any sample that exceeds this latency limit is treated as a failure. All experiments are conducted on the NVIDIA H20-96G GPGPU. We use Qwen2.5-VL-7B-Instruct with Full-Attention as the primary baseline, and compare the performance of Full-Attention and PEVLM under identical latency constraints. To further reflect deployment in low-resource environments, where smaller models are commonly used to meet latency budgets, we also include comparisons with Qwen2.5-VL-3B-Instruct, a smaller variant Qwen2.5-VL-7B-Instruct.
@@
As shown in Table 2, the accuracy of Full-Attention (7B) drops sharply when the latency constraint is reduced from 40s to 30s, and the smaller Full-Attention (3B) model also shows a significant decline when the limit is further tightened to 20s. In contrast, PEVLM maintains consistently high accuracy even under the strictest 20s constraint. Notably, while the larger sink size (sys+16f) achieves the best accuracy under relaxed constraints ($\geq$30s), the smaller configuration (sys+2f) performs better under the 20s limit due to its lower latency overhead.
In general, PEVLM significantly improves accuracy in latency-constrained scenarios. Moreover, to maximize performance across diverse hardware and deployment conditions, the optimal configuration (e.g., sink/block size) should be adaptively tuned.
@@
Sequential Position Embeddin
We further analyze the component with the greatest impact on accuracy: the preservation of sequential position embeddings. As discussed before, adopting sequential position embeddings instead of reusing position embeddings across context blocks produces attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s expected behavior. As shown in Table 4, the use of sequential position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA. In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution (as shown in Figure 9 in the Appendix) remains relatively stable even when using parallel encoding.
@@
PEVLM 58.34% 59.01%
Sink & Context Block Size
As analyzed in the PEVLM section, the inference latency of PEVLM grows quadratically with the sizes of both the sink and context blocks, which aligns with the empirical results shown in Figure 4 and Figure 5. We further investigate how these sizes affect model accuracy.
As noted in the Observations section, including the initial frames of the video in the sink is crucial. This is validated in Figure 6, where incorporating early frames into the sink leads to noticeable accuracy improvements across all evaluated models.
@@
Figure 6 also shows that model accuracy generally increases with sink size. However, due to the quadratic latency growth, excessively large sink sizes are not ideal for deployment. A simple pre-deployment test may be necessary to determine the optimal configuration. For the three models in Figure 6, a sink size between 14 and 18 already achieves strong performance. When latency constraints are considered, this leads to a trade-off between accuracy and efficiency. For example, as shown in Table 2, under a 20-second latency budget, using only [SYS] + 2 frames as the Sink Block yields higher accuracy than [SYS] + 16 frames.
@@
We also examine the effect of context block size on model accuracy. As shown in Figure 7, increasing the size of context blocks leads to consistent accuracy gains, similar to the sink size. However, Figure 4 reveals that latency also increases quadratically with block size, highlighting the same trade-off between accuracy and computational efficiency.
@@ Conclusion
We introduce PEVLM, a fine-tuning-free parallel encoding strategy that significantly accelerates Vision-Language Models for long-video understanding. By preserving sequential position embeddings and leveraging a shared Sink Block, PEVLM aligns attention behavior with Full-Attention while reducing complexity from $O((T\times N)^{2})$ to $O(T\times N)$. It achieves up to 7.47$\times$ attention speedup, 40% end-to-end latency reduction, and strong accuracy gains under tight latency constraints. PEVLM offers a practical and scalable solution for real-world, long-context multimodal tasks.
@@
One additional direction for future exploration is to investigate why parallel encoding methods can surpass Full-Attention in accuracy. Our preliminary analysis suggests that this may be due to the shortened input length in the attention softmax, which improves attention sharpness and reliability in long-context settings. However, since this work focuses on applying and optimizing parallel encoding for VLMs, we leave a deeper theoretical and empirical investigation of this phenomenon for future work. Understanding its root cause may offer new insights into further boosting model performance on long contexts.
Appendix A. Benchmark Details
@@
Appendix B. Attention Weights Distribution of LLaVA-Video and LongVILA